OpenStudy (anonymous):

CAL 3 problem: Working on a divergence theorem question. I don't understand how the answer manual does something. The problem goes... vectors are i j k F= [sqrt(x^2 + y^2 + z^2)] (xi + yj + zk) over the region D: 1<= x^2 + y^2 + z^2 <=2. i get divergence theorem (i think) but the way I went about it was to go through spherical coordinates and let rho=sqrt(x^2 + y^2 + z^2) ->F=rho(x)i + rho(y)j + rho(z)k. then i find the partials wrt x,y,z and come to my roadblock. I get (x^2/rho), same for y and z but my answer manual comes to (x^2/rho) + rho. I don't know where the + rho comes from.

OpenStudy (experimentx):

\( F= \sqrt{x^2 + y^2 + z^2} (xi + yj + zk) \) over the region D: \( 1<= x^2 + y^2 + z^2 <=2 \)

OpenStudy (anonymous):

yes thats correct

OpenStudy (experimentx):

are you trying to take divergence in spherical coordinate system??

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

http://mathworld.wolfram.com/images/equations/SphericalCoordinates/NumberedEquation7.gif this is going to be ugly!!

OpenStudy (anonymous):

even if there is another way to look at this problem, I would like to know how the answer manual gets these partials

OpenStudy (anonymous):

haha i hope not

OpenStudy (experimentx):

It's a concentric sphere

OpenStudy (anonymous):

k. my problem isnt with integrals or divergence theorem. I just need to know where the + rho in the partial derivatives comes from

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=divergence+ [sqrt%28x^2+%2B+y^2+%2B+z^2%29x%2C+sqrt%28x^2+%2B+y^2+%2B+z^2%29y%2C+sqrt%28x^2+%2B+y^2+%2B+z^2%29z] The calculated divergence ... by wolfram alpha

OpenStudy (anonymous):

lol coding fail

OpenStudy (experimentx):

\( \rho \) is the radius of the shpere \[ \rho = \sqrt{x^2 + y^2 + z^2 }\]

OpenStudy (experimentx):

and also what are you trying to calculate??? net flux through the volume??

OpenStudy (anonymous):

flux through volume, ya

OpenStudy (experimentx):

The best way to do is \(\sqrt{x^2 + y^2 + z^2} (xi + yj + zk) = |r|\vec r\) You can calculate \( \nabla. |r|\vec r = 4 |r|\)

OpenStudy (anonymous):

hmm ok i see that, thanks, but i know how to do divergence problems. i just want to know about the + rho part of the partials

OpenStudy (experimentx):

or try changing \[ dV = r^2 \sin\phi \;drd\phi d\theta\] and integrate ... well best of luck!!!

OpenStudy (anonymous):

To other viewers: my problem hasn't been solved. experimentx didn't answer my question.